Two bus A&B travel along the road in same direction from the same starting place .Bus A maintains speed of 120km/h & car B 150km/h but B start 1hour later how many hours will it take for B to overtake A?

Grade:12th pass

1 Answers

Silent Sniper
6 years ago
a=120 b=150 vapp=150-120               distance initlayy=120*1=120 hence we are required to find the time taken to reach a and we know the vapp hence divide it by distance to get time as folows:-
vapp=distance/tie
distance/vapp=time
120/30=time
4hrs=time
this is the answer